# FPC-VLA: A Vision-Language-Action Framework with a Supervisor for Failure Prediction and Correction

### v1

@@ III-B Vision-Language-Action Model
where $\alpha_{t}\in(0,1)$ controls the noise level at timestep $t$, and $x_{0}$ represents the ground truth actions for the next $n$ steps. The model DiT is trained to predict the noise $\epsilon$ added to $x_{0}$, given the noisy input $x_{t}$, the timestep $t$, and the condition $\mathcal{Z}$
$\text{DiT}(x_{t},t,z)\approx\epsilon.$ (2)
III-C Robot Failure Prediction and Correction Dataset
Most existing large-scale VLA datasets [7, 25, 5, 26] only include successful trajectories, while current failure correction datasets [15, 17] are limited in scale and lack predictability. To address this, we propose an automated method for generating failure correction datasets from large-scale RLDS-format data and adapt them for VLM training.
$\mathcal{D}=\{\mathcal{E}_{1},\mathcal{E}_{2},\dots,\mathcal{E}_{N}\},\ \mathcal{E}_{i}=\{s_{1},s_{2},\dots,s_{T}\},$ (3)
where each step $s_{t}$ includes:
  • An RGB image observation $I_{t}\in\mathbb{R}^{H\times W\times 3}$,
  • An action vector $\mathbf{a}_{t}\in\mathbb{R}^{7}$ (6-DOF pose + 1-D gripper),
  • A task-level language instruction $l$.
Discrete grasp events are detected by tracking changes in the gripper state $g_{t}$. The set of change timesteps is
$\mathcal{C}=\{t\mid|g_{t}-g_{t-1}|>\delta_{g}\},$ (4)
where the threshold $\delta_{g}$ is 0.5. For each change timestep $c\in\mathcal{C}$, retain a set of three preceding frames and itself
$\mathcal{R}_{c}=\{c-3,c-2,c-1,c\}\cap[1,T].$ (5)
The total set of retained indices is $\mathcal{R}=\bigcup_{c\in\mathcal{C}}\mathcal{R}_{c}$.
For every retained frame $s_{t}$ with $t\in\mathcal{R}$, we identify the next change timestep $c\in\mathcal{C}$ such that $c\geq t$, and compute the pose difference
$\Delta\mathbf{a}=\mathbf{a}_{c}-\mathbf{a}_{t}.$ (6)
For each $\Delta\mathbf{a}$, we construct a textual description based on movement and rotation direction. For each axis $i\in\{x,y,z\}$, a movement direction and magnitude are defined as
$T_{i}=\left\{\begin{array}[]{l r}\text{None},&|\Delta i|\leq\delta_{\text{trans}}^{\text{small}}\\ \text{Move $D$. Small},&\delta_{\text{trans}}^{\text{small}}<|\Delta i|\leq\delta_{\text{trans}}^{\text{large}}\\ \text{Move $D$. Large},&|\Delta i|>\delta_{\text{trans}}^{\text{large}}\end{array}\right.$.
Around the vertical axis $z$, a rotation direction is defined similarly as
$R=\left\{\begin{array}[]{l r}\text{None},&|\Delta r_{z}|\leq\delta_{\text{rot}}^{\text{small}}\\ \text{Rotate $D$. Small},&\delta_{\text{rot}}^{\text{small}}<|\Delta r_{z}|\leq\delta_{\text{rot}}^{\text{large}}\\ \text{Rotate $D$. Large},&|\Delta r_{z}|>\delta_{\text{rot}}^{\text{large}}\end{array}\right$.,
where $\delta_{\text{trans}}^{\text{small}},\delta_{\text{trans}}^{\text{large}},\delta_{\text{rot}}^{\text{small}},\delta_{\text{rot}}^{\text{large}}$ are thresholds that limit the magnitude of movement or rotation, $\Delta i=i_{t+1}-i_{t}$ represents the position change along the $i\text{-th}$ axis, and $D$ is the direction. Table I illustrates the correspondence between variables and directions using the BridgeV2 [25] dataset as an example.
TABLE I: Mapping of variables and directions in BridgeV2 dataset
Axis Variable Sign Direction $D$
X $\Delta x$ + forward
- backward
Y $\Delta y$ + left
- right
Z $\Delta z$ + up
- down
Around Z $\Delta r_{z}$ + counterclockwise
- clockwise
Given the retained frame $s_{t}$, a standard prompt $\mathcal{P}_{t}$ consists of three parts: task requirements, answer limitations, and example samples. An example is shown in Table II.
TABLE II: Example of failure prediction and correction prompts
Part Content
III-D A VLM-Based Supervisor for Robot Failure Prediction and Correction
Existing robotic manipulation systems based on VLA models often directly execute predicted commands without assessing their contextual appropriateness. These systems typically lack the ability to predict the consequences of actions or to autonomously correct erroneous decisions. To address this limitation, we propose a failure prediction and correction supervisor that evaluates the suitability of gripper actions before execution.
Let the predicted action at timestep $t$ be denoted as:
$\hat{\mathbf{a}}_{t}=[x_{t},y_{t},z_{t},r_{x,t},r_{y,t},r_{z,t},g_{t}],$ (9)
where $(x,y,z)\in\mathbb{R}^{3}$ is the desired end-effector position delta. $(r_{x},r_{y},r_{z})\in\mathbb{R}^{3}$ is the rotation delta. $g$ is the gripper command (0 for close, $1$ for open). Here, $\hat{\mathbf{a}}_{t}$ comes from the output of the action fusion module in Section III-E.
$\left|g_{t}-g_{t-1}\right|>\delta_{g}.$ (10)
When the gripper state $g_{t}$ meets the condition in (10), the prompt $\mathcal{P}_{t}$ and the current image $I_{t}$ are input into a vision language model $\mathcal{M}$
$R_{t}=\mathcal{M}(I_{t},P_{t}),$ (11)
If the response $R_{t}$ is “Yes.”, it indicates that $\mathcal{M}$ approves the decision made by VLA regarding the opening or closing of the gripper, and the action will continue. If $R_{t}$ begins with “No”, the action $\hat{\mathbf{a}}_{t}$ needs to be adjusted according to the modification suggestions provided by $\mathcal{M}$. In the latter case, we extract structured directives from the text, including translation correction $\Delta\mathbf{p}\in\mathbb{R}^{3}$ and rotation correction $\Delta r_{z}\in\mathbb{R}$, based on predefined directional mappings. Each directional token is paired with a discrete magnitude indicator (small or large), which is mapped to a scalar step size ($\delta_{\text{trans}}^{\text{small}}$ or $\delta_{\text{trans}}^{\text{large}}$, $\delta_{\text{rot}}^{\text{small}}$ or $\delta_{\text{rot}}^{\text{large}}$). Based on the language-guided correction, we generate a refined action
$\mathbf{a}^{\prime}_{t}=\hat{\mathbf{a}}_{t}+[\Delta x,\Delta y,\Delta z,0,0,\Delta r_{z},0],$ (12)
leaving unaffected components unchanged. The refined action $\mathbf{a}^{\prime}_{t}$ is then executed within the environment.
III-E Similarity-Guided Action Fusion Module
@@
To obtain a robust and smooth action output and avoid waste in action prediction, we employ an action fusion strategy. At each timestep $t$, the output action $\hat{\mathbf{a}}_{t}$ is not solely determined by the most recent policy output, but rather by fusing a set of historical predictions, each made from earlier observations. This strategy fully takes into account the multimodality and temporal properties of actions, as well as the differences in pose and gripper states.
@@
IV Experiment
TABLE III: Comparison results on SIMPLER with WidowX
Method
Fine-tuned
(in-domain)
Put spoon on towel
Put carrot on plate
Stack green block
on yellow block
Put eggplant
on yellow basket
Avg. Grasp Avg. Success
Grasp Success Grasp Success Grasp Success Grasp Success
RT-1-X [7] 16.7 0.0 20.8 4.2 8.3 0.0 0.0 0.0 11.5 1.1
Octo-Base [8] 34.7 12.5 52.8 8.3 31.9 8.3 66.7 43.1 46.5 16.0
Octo-Small [8] 77.8 47.2 27.8 9.7 40.3 4.2 87.5 56.9 58.4 29.5
OpenVLA [9] 4.1 0.0 33.3 0.0 12.5 0.0 8.3 4.1 14.6 1.0
RoboVLMs [12] 70.8 45.8 33.3 20.8 54.2 4.2 91.7 79.2 62.5 37.5
CogACT [13] 71.7 50.8 15.0 67.5 51.8
SpatialVLA [10]
25.0 20.8 41.7 20.8 58.3 25.0 79.2 70.8 51.2 34.4
SpatialVLA [10]
20.8 16.7 29.2 25.0 62.5 29.2 100.0 100.0 53.1 42.7
ThinkAct [27] 58.3 37.5 8.7 70.8 43.8
FPC-VLA 79.2 79.2 58.3 58.3 70.8 45.8 79.2 75.0 71.9 64.6
TABLE IV: Comparison results on SIMPLER with Google robot

### v2

@@ 3.2 Vision-Language-Action Model
where $\bar{\alpha}_{t}$ is the cumulative product of the noise schedule $\alpha_{t}$ and controls the noise level at timestep $t$, and $x_{0}$ represents the ground-truth action sequence for the subsequent $n$ steps. The DiT model is trained to learn the reverse denoising process. Specifically, given the noisy input $x_{t}$, the timestep $t$, and the conditioning variable $\mathcal{Z}$, the model is tasked to predict the noise component $\epsilon$ that was added to $x_{0}$
$\text{DiT}(x_{t},t,\mathcal{Z})\approx\epsilon.$ (4)
The architecture of the DiT block is pivotal to its performance. The input action sequence $x_{t}$ is first patchified and projected into a sequence of embeddings. The timestep $t$ and the conditioning variable $\mathcal{Z}$ are incorporated into the network through Adaptive Layer Norm (AdaLN), where the parameters for layer normalization are dynamically generated from the embeddings of $t$ and $\mathcal{Z}$. The core of the block consists of Multi-Head Self-Attention and Feed-Forward Network layers, which enable the model to effectively capture the global dependencies that are crucial for coherent multi-step action generation.
During inference, the process starts from a random noise vector $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The trained DiT model then iteratively denoises this vector over $T$ steps, conditioned on $\mathcal{Z}$, to produce a plausible and context-aware action sequence $x_{0}$.
3.3 Dual-Stream Action Fusion Module
@@
To obtain a robust and smooth action output and avoid waste in action prediction, we employ a dual-stream action fusion strategy. At each timestep $t$, the output action $\hat{\mathbf{a}}_{t}$ is not solely determined by the most recent policy output, but rather by fusing a set of historical predictions, each made from earlier observations. This strategy fully takes into account the multimodality and temporal properties of actions, as well as the differences in pose and gripper states.
@@
Algorithm 1 Dual-Stream Action Fusion Strategy
$\text{VLA}(\cdot)$ , $\{\mathcal{O}_{t-k}\mid k=0,1,\dots,N-1\}$ , $\lambda>0$, $\alpha>0$, $\beta\in[0,1)$
Fused action: $\hat{\mathbf{a}}_{t}=[\hat{\mathbf{p}}_{t},\hat{g}_{t}]\in\mathbb{R}^{7}$
Initialize $\mathcal{A}_{t}\leftarrow\emptyset$, $\mathcal{P}_{t}\leftarrow\emptyset$, $\mathcal{G}_{t}\leftarrow\emptyset$
for $k=0$ to N1N-1 do
  Predict action using $\mathcal{O}_{t-k}$: $\mathbf{a}_{t}^{(k)}\leftarrow\text{VLA}(\mathcal{O}_{t-k})$
  Decompose action: $\mathbf{p}_{t}^{(k)},g_{t}^{(k)}]\leftarrow\mathbf{a}_{t}^{(k)}$
  $\mathcal{A}_{t}\leftarrow\mathcal{A}_{t}\cup\{\mathbf{a}_{t}^{(k)}\}$, $\mathcal{P}_{t}\leftarrow\mathcal{P}_{t}\cup\{\mathbf{p}_{t}^{(k)}\}$, $\mathcal{G}_{t}\leftarrow\mathcal{G}_{t}\cup\{g_{t}^{(k)}\}$
end for
for $k=0$ to N1N-1 do
  $\text{sim}^{(k)}\leftarrow\dfrac{\mathbf{p}_{t}^{(k)}\cdot\mathbf{p}_{t}^{(0)}}{\|\mathbf{p}_{t}^{(k)}\|\cdot\|\mathbf{p}_{t}^{(0)}\|+\varepsilon}$ $\triangleright$ Cosine Similarity
  Apply sigmoid transformation: $s^{(k)}\leftarrow\sigma(\alpha\cdot\text{sim}^{(k)})$
end for
for $k=0$ to N1N-1 do
  $d^{(k)}\leftarrow\exp(-\lambda\cdot k)$ $\triangleright$ Time Decay
end for
$S\leftarrow 0$
for $k=0$ to N1N-1 do
  $S\leftarrow S+\max(0,(1-\beta)\cdot s^{(k)}\cdot d^{(k)})$ $\triangleright$ Regularization
end for
for $k=0$ to N1N-1 do
  $w^{(k)}\leftarrow\dfrac{\max(0,(1-\beta)\cdot s^{(k)}\cdot d^{(k)})}{S}$
end for
$\hat{\mathbf{p}}_{t}\leftarrow\mathbf{0}\in\mathbb{R}^{6}$
for $k=0$ to N1N-1 do
  $\hat{\mathbf{p}}_{t}\leftarrow\hat{\mathbf{p}}_{t}+w^{(k)}\cdot\mathbf{p}_{t}^{(k)}$ $\triangleright$ Pose Fusion
end for
$D\leftarrow\sum_{j=0}^{N-1}d^{(j)}$
for $k=0$ to N1N-1 do
  $\tilde{d}^{(k)}\leftarrow\dfrac{d^{(k)}}{D}$
end for
$\hat{g}_{t}\leftarrow 0$
for $k=0$ to N1N-1 do
  $\hat{g}_{t}\leftarrow\hat{g}_{t}+\tilde{d}^{(k)}\cdot g_{t}^{(k)}$ $\triangleright$ Gripper State Fusion
end for
$\hat{\mathbf{a}}_{t}\leftarrow[\hat{\mathbf{p}}_{t},\hat{g}_{t}]$
return $\hat{\mathbf{a}}_{t}$
@@
3.4 Robot Failure Prediction and Correction Dataset
Most existing large-scale VLA datasets [o2024open, walke2023bridgedata, brohan2022rt, liu2023libero] only include successful trajectories, while current failure correction datasets [luo2025roboreflect, lu2025robofac] are limited in scale and lack predictability. To address this, we propose an automated method for generating failure correction datasets from large-scale RLDS-format data and adapt them for VLM training. In practical experiments, in addition to synthetic data, simulation platforms such as MuJoCo and teleoperation methods with real robots are used to construct failure correction data through annotation.
$\mathcal{D}=\{\mathcal{E}_{1},\mathcal{E}_{2},\dots,\mathcal{E}_{N}\},\ \mathcal{E}_{i}=\{s_{1},s_{2},\dots,s_{T}\}.$ (13)
At each timestep $t$, the step $s_{t}$ encapsulates the full state of the system, defined by the tuple
$s_{t}=(I_{t},\mathbf{a}_{t},l),$ (14)
where each step $s_{t}$ includes:
  • 1.
    An RGB image observation $I_{t}\in\mathbb{R}^{H\times W\times 3}$,
  • 2.
    An action vector $\mathbf{a}_{t}\in\mathbb{R}^{7}$, typically comprising a 6-DOF end-effector pose change and a 1-D gripper state.
  • 3.
    A task-level language instruction $l$.
We posit that grasp and release events are pivotal moments in manipulation tasks, often signifying the transition between task phases. Let $g_{t}\in[0,1]$ be the normalized gripper state at time $t$ (e.g., 0 for open, 1 for closed). The set of all gripper state change timesteps $\mathcal{C}$ is defined as
$\mathcal{C}=\left\{t\in\{2,\dots,T\}\ \middle|\ |g_{t}-g_{t-1}|>\delta_{g}\right\},$ (15)
where $\delta_{g}$ is a sensitivity threshold, empirically set to $\delta_{g}=0.5$. To provide sufficient contextual information for a VLM to understand the scene dynamics leading up to a change, we retain a window of frames around each event. For a given change timestep $c\in\mathcal{C}$, the set of retained indices is
$\mathcal{R}_{c}=\{c-k,\dots,c-1,c\}\cap[1,T],$
where $k$ is the context window size. In our implementation, $k=3$. The union of all these sets forms the complete set of frames used for dataset generation
$\mathcal{R}=\bigcup_{c\in\mathcal{C}}\mathcal{R}_{c}.$
For every retained frame $s_{t}$ where $t\in\mathcal{R}$, we generate a corrective action target. This is achieved by identifying the next gripper change event that the robot must execute. Formally, we find:
$c^{*}=\min\{c\in\mathcal{C}\mid c\geq t\}.$
The corrective action vector $\Delta\mathbf{a}$ is then computed as the difference between the action at this future event and the current action:
$\Delta\mathbf{a}=\mathbf{a}_{c^{*}}-\mathbf{a}_{t}.$
This vector $\Delta\mathbf{a}$ encodes the necessary pose adjustment to progress towards the next critical phase of the task.
As indicated by the pre-training process of VLA [zheng2024tracevla, li2024towards, li2024cogact, qu2025spatialvla], directly obtaining continuous action corrections requires large-scale pre-training data and high training costs. Therefore, we map the continuous correction values into discrete natural language signals, thereby fulfilling the task requirements with only LoRA fine-tuning of the VLM. To make the corrective action comprehensible to a VLM, we map the continuous vector $\Delta\mathbf{a}$ into a structured natural language description. We decompose $\Delta\mathbf{a}$ into its translational ($\Delta x,\Delta y,\Delta z$) and rotational (primarily around the vertical axis, $\Delta r_{z}$) components.
Let $D_{i}$ be the direction for axis $i$ (e.g., ‘left’, ‘right’, ‘up’, ‘down’), determined by the sign of $\Delta i$. The magnitude-based textual template for translation along axis $i$ is generated by the function:
$\text{Text}_{\text{trans}}(i)=\begin{cases}\text{None},&|\Delta i|\leq\delta_{\text{small}}\\ \text{Move $D_{i}$. Small},&\delta_{\text{small}}<|\Delta i|\leq\delta_{\text{large}}\\ \text{Move $D_{i}$. Large},&|\Delta i|>\delta_{\text{large}}\end{cases}$
Similarly, the rotation around the $z$-axis is described by:
$\text{Text}_{\text{rot}}(z)=\begin{cases}\text{None},&|\Delta r_{z}|\leq\delta_{\text{small}}\\ \text{Rotate $D_{z}$. Small},&\delta_{\text{small}}<|\Delta r_{z}|\leq\delta_{\text{large}}\\ \text{Rotate $D_{z}$. Large},&|\Delta r_{z}|>\delta_{\text{large}}\end{cases}$
where $D_{z}$ is either ‘clockwise’ or ‘counter-clockwise’. Since rotations around the xy-axis are difficult to translate into natural language and VLMs have a poor understanding of them, here we only consider the most frequently occurring clockwise and counterclockwise rotation issues. In practice, $\delta_{\text{small}}=0.01$, $\delta_{\text{large}}=0.1$. The final instruction is a concatenation of all non-”None” components from $\text{Text}_{\text{trans}}(x)$, $\text{Text}_{\text{trans}}(y)$, $\text{Text}_{\text{trans}}(z)$, and $\text{Text}_{\text{rot}}(z)$. The specific mapping between variables and directional language is illustrated in Table 1. Due to the consistent coordinate system in VLA training data such as [walke2023bridgedata, brohan2022rt], the prompts indicate that all directions are defined relative to the robot’s base frame, rather than from the camera’s perspective.
Table 1: Mapping of variables and directions in BridgeV2 dataset
Axis Variable Sign Direction $D$
X $\Delta x$ + forward
- backward
Y $\Delta y$ + left
- right
Z $\Delta z$ + up
- down
Around Z $\Delta r_{z}$ + counterclockwise
- clockwise
Given the retained frame $s_{t}$, a standard prompt $\mathcal{P}_{t}$ consists of three parts: task requirements, answer limitations, and example samples. An example is shown in Table 2.
Table 2: Example of failure prediction and correction prompts
Part Content
3.5 A VLM-Based Supervisor for Robot Failure Prediction and Correction
Existing robotic manipulation systems based on VLA models often directly execute predicted commands without assessing their contextual appropriateness. These systems typically lack the ability to predict the consequences of actions or to autonomously correct erroneous decisions. To address this limitation, we propose a supervisor for failure prediction and correction, which evaluates the suitability of gripper actions before execution. The algorithm workflow is shown in Algorithm 2.
Algorithm 2 Supervisor for Failure Prediction and Correction
1:Vision-Language Model $\mathcal{M}$, Prompt Template $\mathcal{P}$, Threshold $\delta_{g}$
2:Initialize $g_{t-1}\leftarrow 0$
3:for each timestep $t=1,2,\dots$ do
4:  Receive $\hat{\mathbf{a}}_{t}=[\Delta x_{t},\Delta y_{t},\Delta z_{t},\Delta r_{x,t},\Delta r_{y,t},\Delta r_{z,t},g_{t}]^{\top}$ from action fusion module
5:  Receive current image observation $I_{t}$
6:  if $|g_{t}-g_{t-1}|>\delta_{g}$ then
7:    $R_{t}\leftarrow\mathcal{M}(I_{t},\mathcal{P}_{t})$
8:    if $R_{t}$ begins with ”No” then
9:     Extract $\Delta\mathbf{p}=[\Delta x,\Delta y,\Delta z]$ and $\Delta r_{z}$ from $R_{t}$
10:     Map directional tokens to step sizes: $\delta_{\text{small}}$, $\delta_{\text{large}}$
11:     $\mathbf{a}^{\prime}_{t}\leftarrow\hat{\mathbf{a}}_{t}+[\Delta x,\Delta y,\Delta z,0,0,\Delta r_{z},0]$
12:     Execute $\mathbf{a}^{\prime}_{t}$
13:    else
14:     Execute $\hat{\mathbf{a}}_{t}$
15:    end if
16:  else
17:    Execute $\hat{\mathbf{a}}_{t}$
18:  end if
19:  $g_{t-1}\leftarrow g_{t}$
20:end for
Let the raw action command predicted by the primary VLA policy at timestep $t$ be parameterized as a 7D vector:
$\hat{\mathbf{a}}_{t}=[\Delta x_{t},\Delta y_{t},\Delta z_{t},\Delta r_{x,t},\Delta r_{y,t},\Delta r_{z,t},g_{t}]^{\top},$ (16)
where $(\Delta x,\Delta y,\Delta z)\in\mathbb{R}^{3}$ denotes the desired end-effector positional displacement in the robot’s base frame. $(\Delta r_{x},\Delta r_{y},\Delta r_{z})\in\mathbb{R}^{3}$ represents the incremental rotation, and $g_{t}\in\{0,1\}$ is the binary gripper state command. This action vector $\hat{\mathbf{a}}_{t}$ is the direct output from the action fusion module detailed in Section 3.3.
$\left|g_{t}-g_{t-1}\right|>\delta_{g}.$ (17)
A transition satisfying (17) signifies an intended change in the gripper’s state, which is a high-stakes decision requiring contextual validation. When the gripper state $g_{t}$ meets the condition in (17), the prompt $\mathcal{P}_{t}$ and the current image $I_{t}$ are input into a vision language model $\mathcal{M}$
$R_{t}=\mathcal{M}(I_{t},P_{t}),$ (18)
If the response $R_{t}$ is “Yes.”, it indicates that $\mathcal{M}$ approves the decision made by VLA regarding the opening or closing of the gripper, and the original command $\hat{\mathbf{a}}_{t}$ is executed without modification. If $R_{t}$ begins with “No”, the action $\hat{\mathbf{a}}_{t}$ needs to be adjusted according to the modification suggestions provided by $\mathcal{M}$. In the latter case, we extract structured directives from the text, including translation correction $\Delta\mathbf{p}\in\mathbb{R}^{3}$ and rotation correction $\Delta r_{z}\in\mathbb{R}$, based on predefined directional mappings. Each directional token is paired with a discrete magnitude indicator (small or large), which is mapped to a scalar step size.The refined action $\mathbf{a}^{\prime}_{t}$ is subsequently generated by applying the parsed correction to the original command, selectively overriding only the components for which a correction was specified
$\mathbf{a}^{\prime}_{t}=\hat{\mathbf{a}}_{t}+[\Delta x,\Delta y,\Delta z,0,0,\Delta r_{z},0].$ (19)
The refined action $\mathbf{a}^{\prime}_{t}$ is then executed within the environment.
4 Experiment
Table 3: Comparison results on SIMPLER with WidowX
Method Fine-tuned (in-domain) Put spoon on towel Put carrot on plate Stack green block on yellow block Put eggplant on yellow basket Avg. Grasp Avg. Task Grasp Task Grasp Task Grasp Task Grasp Task RT-1-X [o2024open] 16.7 0.0 20.8 4.2 8.3 0.0 0.0 0.0 11.5 1.1 Octo-Base [team2024octo] 34.7 12.5 52.8 8.3 31.9 8.3 66.7 43.1 46.5 16.0 Octo-Small [team2024octo] 77.8 47.2 27.8 9.7 40.3 4.2 87.5 56.9 58.4 29.5 OpenVLA [kim2024openvla] 4.1 0.0 33.3 0.0 12.5 0.0 8.3 4.1 14.6 1.0 RoboVLMs [li2024towards] 70.8 45.8 33.3 20.8 54.2 4.2 91.7 79.2 62.5 37.5 CogACT [li2024cogact] 71.7 50.8 15.0 67.5 51.8 SpatialVLA [qu2025spatialvla] 25.0 20.8 41.7 20.8 58.3 25.0 79.2 70.8 51.2 34.4 SpatialVLA [qu2025spatialvla] 20.8 16.7 29.2 25.0 62.5 29.2 100.0 100.0 53.1 42.7 ThinkAct [thinkact] 58.3 37.5 8.7 70.8 43.8 FPC-VLA 79.2 79.2 58.3 58.3 70.8 45.8 79.2 75.0 71.9 64.6
Table 4: Comparison results on SIMPLER with Google robot